OpenStudy (anonymous):

Which graph describes the open sentence 1 ≤ n + 4 ≤ 8?

OpenStudy (nincompoop):

which which which which? where are the choices to pick from? WHICH ENTAILS A LIST OF CHOICE

OpenStudy (anonymous):

subtract 4 from each inequality

OpenStudy (anonymous):

1 <= n + 4 <= 8 -- subtract 4 1 - 4 <= n + 4 - 4 <= 8 - 4 -- combine like terms -3 <= n <= 4